### v1

# Computational considerations for the prediction of airfoil Stall Flutter

Abstract
This paper presents a comprehensive numerical investigation of a NACA0012 undergoing Stall Flutter Limit Cycle Oscillations (LCO) across distinct fluid dynamics regimes. It accurately models Small Amplitude Oscillations (SAO) in the transitional Reynolds regime and Large Amplitude Oscillations (LAO) in the moderate regime, observed in different experimental campaigns. The SAO analysis servs as a verification of the computational framework against established numerical benchmarks. Crucially, the LAO simulations represent the first documented prediction across the full experimental velocity range correlated against available measured data, addressing a significant literature gap. The predictions fidelity relies on rigorous computational criteria defined through a detailed sensitivity analysis. This demonstrated numerical requirements significantly more demanding than those typically employed for computing static polars or simulating dynamic pitching motion of rigid airfoils, underscoring the severity of the aeroelastic problem. Overall, the results show strong qualitative agreement with experimental observations, successfully reproducing key dynamic stall mechanics and bifurcation phenomena. Quantitatively, however, the simulation systematically over–predicts the critical onset velocity and under–predicts the LCO amplitudes, a discrepancy attributed to reduced aerodynamic excitation peaks.
Keywords Fluid Structure Interaction (FSI) $\cdot$ Stall Flutter $\cdot$ Limit Cycle Oscillations (LCO) $\cdot$ NACA0012 $\cdot$ Dynamic Stall $\cdot$ 2D URANS
1 Introduction
Stall flutter is a non–linear aeroelastic instability phenomenon resulting from the coupling between the flexibility of an aerodynamic structure (a wing) and the flow, in cases fundamentally driven by dynamic flow separation (dynamic stall). The resulting structural response ends up in a change in the pitch angle of the wing, which in turn may give rise to periodic complex flow separation and perhaps re–attachment on both sides of the lifting surface. This interaction between the fluid and the structure (Fluid Structure Interaction – FSI) creates self–excited oscillations and allows for the continued energy transfer from the free stream to the wing. Depending on the correlation between the aerodynamic energy input and the structural energy dissipation, the system’s dynamic behavior varies. Classical bending–torsion flutter involves the interaction between two (or more) structural modes, leading to self–excited vibrations of exponentially increasing amplitude that are mainly unrelated to flow separation (at least in the initial phase of the phenomenon) and usually occur in a linear flow regime (see Figure 1). On the other hand, stall flutter is strongly related to non–linear aerodynamic excitation phenomena arising in stall and post–stall angles of attack, leading in many cases to self–excited single degree of freedom (pitching) oscillations of limited amplitude that are usually referred to as Limit Cycle Oscillations (LCO). LCO represent a state of dynamic equilibrium where the time–averaged energy absorbed from the fluid over one cycle is balanced by the energy dissipated by structural damping. This instability is commonly encountered on similar aerodynamic shapes such as wings or rotating blades in many different applications, including helicopter rotor blades ham1966torsional, wind turbine blades ericsson2000role, turbomachinery blades mendelson1949aerodynamic, aircraft tails victory1943flutter and highly flexible wings noll2004investigation.
Although initial experimental work on stall flutter dates back to the 1940s victory1943flutter, much of the related research has historically focused on the aerodynamic phenomenon of dynamic stall in the second half of the $20^{th}$ century halfman1951evaluation; ericsson1980dynamic; mccroskey1981phenomenon; rivera1992naca. Nevertheless, stall flutter is an aeroelastic phenomenon resulting in LCO introduced by complex and non–linear dynamic flow phenomena such as dynamic stall. Dynamic stall is a distinct unsteady aerodynamic phenomenon, in which an abrupt loss of aerodynamic loads occurs due to flow separation over a lifting surface undergoing unsteady motion that may not necessarily originate in aeroelastic responses. Yet, in stall flutter oscillations, dynamic stall serves as the main excitation factor that stabilizes the limited amplitude oscillations. Over recent years, various types of low–speed stall flutter oscillations have been observed experimentally, leading to several classifications, typically based on the dominant degree of freedom (e.g. pitch-dominated, plunge-dominated). Within the pitch-dominated classification, distinct behaviors are observed:
  • Small Amplitude Oscillations (SAO) with symmetric pitch response in the range between $0^{\circ}$ and $10^{\circ}$ around $0^{\circ}$ mean value were first systematically observed and reported through wind tunnel experiments in the transitional Reynolds number regime $(4.5*10^{4}\leq Re_{c}\leq 1.3*10^{5})$ poirel2008self. This type of SAO arises because laminar boundary layer separation near the trailing edge region initiates the loss of dynamic stability at $0^{\circ}$ pitch angle, while the subsequent presence of a Laminar Separation Bubble (LSB) stabilizes the fluttering motion into LCO. This transitional instability regime has since been thoroughly investigated using various Computational Fluid Dynamics (CFD) tools. Early computational efforts to numerically investigate the flow patterns generated around the oscillating wing and the resulting aerodynamics loads utilised Large Eddy Simulations (LES) relying on imposed motion poirel2010aerodynamics, while later analyses successfully employed fully unsteady aeroelastic simulations based on a 2D Unsteady Reynolds Averaged Navier Stokes (URANS) approach, effectively reproducing the self–sustained pitch oscillations across the relevant Reynolds number range poirel2011computational; abbas2021aeroelastic. These numerical predictions are utilised in the present work for the verification of our comprehensive computational framework, allowing for crucial cross–reference against external data of comparable fidelity.
  • Large Amplitude Oscillations (LAO) with symmetric pitch response in the range between $10^{\circ}$ and $40^{\circ}$ that are observed in a higher Reynolds number range $(2.5*10^{5}\leq Re_{c}\leq 5*10^{5})$ and are linked to the generation and shedding of a strong Leading Edge Vortex (LEV), similar to those observed during dynamic stall li2007experimental; dimitriadis2009bifurcation. In the same experimental campaign, asymmetric LCO of smaller amplitude (up to $20^{\circ}$) around the static equilibrium angle have been also observed at higher Reynolds numbers (up to $5*10^{5}\leq Re_{c}\leq 6*10^{5}$), resulting from the interaction between the stall flutter mechanism and the system’s static divergence. Contrary to SAO cases, accurate CFD predictions reproducing these large amplitude LCO experiments are relatively scarce in the literature yabili2012unsteady. At least to the authors’ knowledge, the successful computational replication of these large amplitude stall flutter LCO across the entire experimental Reynolds number range has not yet been thoroughly reported, highlighting a significant literature gap that this paper aims to address.
[…]
The structure of the paper is as follows. Section 2 describes the experimental setups used for validation, focusing on small and large amplitude oscillations. Section 3 details the computational methodology, including the aerodynamic and structural models used, the aeroelastic coupling approach, and the comprehensive sensitivity analysis performed for grid and time–step dependency. Section 4 presents the numerical results, comparing predicted LCO characteristics and flow physics against experimental measurements and computational predictions (wherever available) for both small and large amplitude stall flutter cases. Finally, Section 5 concludes the paper by summarizing the key findings and contributions.
Figure 1: Schematic diagram of different flutter types.
2 Experimental set–up
The following sections detail the experimental configurations utilised for the validation the computational framework presented in this paper, focusing on the distinct stall flutter regimes: Small Amplitude Oscillations (SAO) and Large Amplitude Oscillations (LAO). Sections and 4.2 provide specific details on the physical set–ups and structural parameters of these experiments, while Section 2.3, offers a general overview of the numerical approach.
2.1 Small Amplitude Oscillations
The validation case for SAO is based on experiments performed by Poirel, Harris and Benaissa poirel2008self in the closed circuit low–speed wind tunnel of the Royal Military College of Canada (RMC). This study focused on aeroelastic LCO occurring in the narrow transitional Reynolds number regime, specifically $4.5*10^{4}\leq Re_{c}\leq 1.3*10^{5}$. The model utilised a NACA0012 airfoil section (chord c=0.156mc=0.156\;m, span b=0.61mb=0.61\;m) that was structurally constrained to rotate in pure pitch, with plunging motion held fixed due to infinite stiffness, and was mounted with end–plates in order to minimize $3D$ effects. The apparatus possessed a mass moment of inertia $I_{s}=0.00135\;kgm^{2}$ and a structural stiffness $K_{s}=0.30\;Nm/rad$ and in the baseline configuration, the elastic axis (pitching axis) was located at $X_{ea}=0.18c$ aft of the leading edge. The system exhibited a viscous structural damping coefficient $D_{s}=0.002\;Nms$, resulting in a structural damping ratio $\zeta=0.05$ during free decay tests in still air. The self–sustained oscillations were typically observed for airspeeds between $5$ and 12m/s12\;m/s, corresponding to the transitional Reynolds range. The baseline value for the free–stream turbulence intensity of the wind tunnel was $T_{u}\approx 0.2\%$. These small amplitude oscillations ($\theta_{max}\leq 5.5^{\circ}$) were the first LCO associated with low Reynolds number effects to be systematically reported. The underlying mechanism is fundamentally linked to laminar boundary layer separation near the trailing edge, initiating the instability and leading to the designation laminar separation flutter. The motion was generally characterized by a well–behaved harmonic response and was measured with rotary potentiometers, whereas the flow periodicity was monitored using hot–wire anemometry in the wake. The aerodynamic moment was calculated as the right hand side of the dynamic equation of motion of the wing, by assuming a pure sinusoidal variation of the pitch angle of the wing with a known frequency and amplitude. A schematic of the experimental set–up is shown in Figure 2.
Figure 2: Schematic of the experimental set–up used to study Small Amplitude Stall Flutter Oscillations. Image adapted from poirel2008self.
2.2 Large Amplitude Oscillations
The LAO experiments used for validation were conducted at the University of Manchester by Dimitriadis and Li li2007experimental; dimitriadis2009bifurcation. Static measurements were carried out in the Project tunnel, which is an open–return subsonic wind tunnel with a working section of 2.16m2.16\;m in length, 1.1m1.1\;m in width and 0.86m0.86\;m in height, whereas dynamic measurements were conducted in the Avro $9x7\;ft^{2}$ closed–return subsonic wind tunnel. The test model featured a rectangular wing with a NACA0012 airfoil section, defined by a b=900mmb=900\;$m$ span and a c=300mmc=300\;$m$ chord. The support mechanism allowed for two degrees of freedom (pitch and plunge). However, the configuration was engineered to exhibit a single–mode dynamic response dominated by pitch motion due to the intentionally high plunge stiffness ($K_{h}=30.5\;N/mm$) compared to the pitch stiffness ($K_{p}=13.1\;Nm/rad$). The pitch axis was set $X_{ea}\approx 0.383c$ aft of the leading edge. The wing had a total weight of m=15kgm=15\;kg and a moment of inertia of $I_{p}=0.31\;kgm^{2}$ about the pitch axis. The system did not apply explicit damping mechanisms. Instead, the observed damping was primarily due to friction in the pitch springs, which led to non–linear structural damping. Analysis of free decay tests at 0, $2.5$ and 3.1m/s3.1\;m/s airspeed showed that the damping of the pitching mechanism was essentially linear at high pitch angles but became significantly non–linear at low pitch angles. The damping in plunge can be assumed to be very large. This campaign captured symmetric LCO with amplitudes reaching up to $40^{\circ}$, typically observed for airspeeds between $12$ and 25m/s25\;m/s that correspond to a moderate Reynolds number range extending from $2.4*10^{5}$ up to $5*10^{5}$. Furthermore, asymmetric LCO were observed at higher airspeeds between $25$ and 30m/s30\;m/s ($5*10^{5}\leq Re_{c}\leq 6*10^{5}$) resulting from the interaction of the flutter mechanism with the static divergence airspeed of the system. No self–sustained oscillations were observed for airspeeds between $3$ and 12m/s12\;m/s ($6*10^{4}\leq Re_{c}\leq 2.4*10^{5}$). The free–stream turbulence intensity in this large wind tunnel was found to be low, typically estimated to be around $T_{u}=0.05\%$. Unsteady pressure data were measured around the mid–span section, whereas the motion of the wing was measured using laser displacement probes. A picture of the experimental set–up is shown in Figure 3.
Figure 3: Image of the experimental set–up used to study Large Amplitude Stall Flutter Oscillations. Image copied from li2007experimental.
2.3 Computational set–up
[…]
Figure 4: Schematic of the computational set–up used to study the Small Amplitude Stall Flutter Oscillations measured in poirel2008self and the Large Amplitude Stall Flutter Oscillations measured in dimitriadis2009bifurcation. The aerodynamic analysis is performed through a 2D URANS CFD framework, whereas for the prediction of the system dynamics response, only the pitch DOF is considered.
3 Computational Methodology
This chapter presents the comprehensive computational methodology employed to simulate the aeroelastic response of the NACA0012 airfoil under stall flutter conditions, as defined by the experimental setups in Section 2. The foundation of this analysis is the in–house CFD solver MaPFlow, which is strongly coupled to an internal 6 DOFs RBD solver for the analysis of FSI problems.
3.1 The aerodynamic model
The core aerodynamic calculations for the aeroelastic simulations are performed using the in–house CFD solver MaPFlow papadakis2014development, which utilises a cell–centered finite volume discretisation scheme and is capable of handling both structured and unstructured computational grids. MaPFlow is built upon the compressible URANS equations described by Equation (1).
$\frac{\partial}{\partial t}\int_{D}\mathbf{U}dD+\oint_{\partial{D}}\left(\mathbf{F_{c}}-\mathbf{F_{v}}\right)dS=\int_{D}\mathbf{Q}dD$ (1)
In Equation (1), $\mathbf{U}$ is the vector of the Conservative Flow Variables:
$\mathbf{U}=\left(\begin{array}[]{ccc}\rho\\ \rho u\\ \rho v\\ \rho w\\ \rho E\end{array}\right)$
where $\rho$ denotes density, $\left(u,v,w\right)$ the three components of the velocity field and $E$ the total energy.
The convective fluxes $\mathbf{F_{c}}$, described by Equation (2), are evaluated using the Roe’s approximate Riemann solver roe1986characteristic described in Equation (3) coupled with the Venkatakrishnan limiter venkatakrishnan1993accuracy.
$\mathbf{F_{c}}=\left(\begin{array}[]{ccc}\rho V\\ \rho uV+n_{x}p\\ \rho vV+n_{y}p\\ \rho wV+n_{z}p\\ \rho(E+\frac{p}{\rho})V\end{array}\right)$ (2)
$\mathbf{F}_{c_{f}}=\frac{1}{2}[\mathbf{F}_{c}(\mathbf{V}_{R})+\mathbf{F}_{c}(\mathbf{V}_{L})-|\mathbf{A_{Roe}}|_{f}(\mathbf{V}_{R}-\mathbf{V}_{L})]$ (3)
The Left and Right states $\left(\mathbf{V}_{L},\mathbf{V}_{R}\right)$ of the primitive variables $\left(\mathbf{V}=\left(\rho,u,v,w,p\right)^{T}\right)$ of the face are computed using a Piecewise Linear Reconstruction (PLR) scheme and $|\mathbf{A_{Roe}}|$ is constructed using the absolute values of the eigenvalues and the right eigenvector matrix $R$:
$|\mathbf{A_{Roe}}|=\mathbf{R}^{-1}|\boldsymbol{\Lambda}|\mathbf{R}$
In case Low Mach Preconditioning is used, the eigenvalues and the eigenvectors of the preconditioned system must be used ($\boldsymbol{\Lambda}_{\Gamma},\mathbf{R}_{\Gamma}$). Hence, $|\mathbf{A_{Roe}}|$ is changed to (according to Asouti_phd):
$|\mathbf{A}_{\Gamma\mathbf{Roe}}|$ $=|\boldsymbol{\Gamma}^{-1}\boldsymbol{\Gamma}\mathbf{A_{Roe}}|$
$\simeq\boldsymbol{\Gamma}^{-1}|\boldsymbol{\Gamma}\mathbf{A_{Roe}}|$
$\simeq\boldsymbol{\Gamma}^{-1}\mathbf{R}^{-1}_{\Gamma}|\boldsymbol{\Lambda}_{\Gamma}|\mathbf{R}_{\Gamma}$
The viscous fluxes $\mathbf{F_{v}}$, described by Equation (4), are calculated using a central $2nd$ order scheme.
$\mathbf{F_{v}}=\left(\begin{array}[]{ccc}0\\ n_{x}\tau_{xx}+n_{y}\tau_{xy}+n_{z}\tau_{xz}\\ n_{x}\tau_{yx}+n_{y}\tau_{yy}+n_{z}\tau_{yz}\\ n_{x}\tau_{zx}+n_{y}\tau_{zy}+n_{z}\tau_{zz}\\ n_{x}\Theta_{x}+n_{y}\Theta_{y}+n_{z}\theta_{z}\end{array}\right)$ (4)
In Equations (2) and (4) $V=\mathbf{u}\cdot\mathbf{n}$, $p$ denotes pressure, $\boldsymbol{\tau}$ is the viscous stress tensor and
$\Theta_{x}$ $=u\tau_{xx}+v\tau_{xy}+w\tau_{xz}$
$\Theta_{y}$ $=u\tau_{yx}+v\tau_{yy}+w\tau_{yz}$
$\Theta_{z}$ $=u\tau_{zx}+v\tau_{zy}+w\tau_{zz}$
The above system is completed with the Equation (5) of state for perfect gases:
$p=(\gamma-1)\rho\left[E-\frac{u^{2}+v^{2}+w^{2}}{2}\right]$ (5)
[…]
MaPFlow can handle both steady and unsteady flows. The unsteady nature of FSI problems necessitates time–true computations using the implicit $2nd$ order backwards differences scheme described in Equation (6), combined with the dual time–stepping technique to facilitate convergence vatsa2010re.
$\frac{d\left(D_{I}\mathbf{U}_{I}\right)}{dt}+\mathbf{R}^{n+1}_{I}=\mathbf{0}$ (6)
where
$\frac{d\left(D_{I}\mathbf{U}_{I}\right)}{dt}=\frac{1}{\Delta t}\left[3/2\left(D_{I}\mathbf{U}_{I}\right)^{n+1}-2\left(D_{I}\mathbf{U}_{I}\right)^{n}+1/2\left(D_{I}\mathbf{U}_{I}\right)^{n-1}\right],$
$\mathbf{R}^{n+1}\approx\mathbf{R}^{n}+\left(\frac{\partial\mathbf{R}}{\partial\mathbf{U}}\right)_{n}\Delta\mathbf{U}^{n},$
$\mathbf{R}_{I}=\left[\sum\limits_{f=1}^{N_{f}}\left(\mathbf{F}_{c_{f}}-\mathbf{F}_{v_{f}}\right)\Delta S_{f}-D\mathbf{Q}\right]_{I}$
and
$\Delta\mathbf{U}^{n}=\mathbf{U}^{n+1}-\mathbf{U}^{n}$
To ensure numerical stability and robustness in the low Mach number flow regime, Eriksson’s Preconditioning Matrix Eriksson1996 is consistently applied to the governing equations, due to its successful use in Asouti2005.
3.2 The structural dynamics model
The structural response of the airfoil is handled internally by the in–house RBD solver of MaPFlow, which is integrated within the numerical framework MaPFlow and is strongly coupled with the aerodynamic system to perform high–fidelity FSI computations. For the validation study presented here, the RBD solver is strictly constrained to model only the pitch DOF. The governing equation of motion for pitch rotation $\theta$ is expressed as a standard second–order system:
$I\ddot{\theta}+C\dot{\theta}+K\theta=M_{aero}$ (7)
, where $I$ is the mass moment of inertia, $C$ the damping coefficient and the $K$ is the stiffness of the torsional spring. $M_{aero}$ is the aerodynamic moment calculated by properly integrating pressure and viscous stresses over the boundary of the airfoil following equation (8):
$M_{aero}$ $=\int_{\partial B}\mathbf{r}\times(-p\mathbf{n}+\boldsymbol{\tau}\cdot\mathbf{n})dS$ (8)
, where the $p$ is the pressure distribution over the lifting surface, $\boldsymbol{\tau}$ is the viscous stress tensor, $\mathbf{n}$ the outwards pointing unit normal vector of the surface, $\mathbf{r}$ the position vector with its origin being at the center of rotation and $\partial B$ is the surface boundary of the rigid body.
[…]
3.3 The aeroelastic coupling
A strong aeroelastic coupling between the fluid domain and the structural dynamics solver is achieved through an iterative solution of the two separate, fluid and structural, problems within every physical time step to ensure convergence of the fluid–structure equilibrium. The overall FSI process involves the following exchange:
  1. 1.
    Given a flow field solution, the CFD solver computes the aerodynamic moment ($M_{aero}$) acting on the airfoil.
  2. 2.
    The aerodynamic moment is transferred to the RBD solver, which integrates the equation of motion (7) to determine the new pitch angle $\theta$, angular velocity $\dot{\theta}$ and angular acceleration $\ddot{\theta}$ of the airfoil.
  3. 3.
    The updated kinematic state (orientation and velocity) is communicated back to the flow solver by properly rotating the computational grid.
  4. 4.
    The iterative loop continues until the calculated motion change with respect to the previous iteration achieves a predetermined tolerance of $\epsilon=10^{-4}$: $\max(|{\theta}^{n+1}-{\theta}^{n}|,|{\dot{\theta}}^{n+1}-{\dot{\theta}}^{n}|,|{\ddot{\theta}}^{n+1}-{\ddot{\theta}}^{n}|)\leq\epsilon$.
The typical workflow of this iterative, strongly coupled procedure is detailed in Figure 5.
It needs to be highlighted that the computational mesh used in this study is rotated rigidly to accommodate the computed pitching motion. This approach is preferred to mesh deformation techniques as it conserves the regularity of the grid, thus reducing the numerical diffusion associated with mesh deformation, and ensures ease of implementation, thus restraining numerical errors. The rigid body rotation is implemented by employing the 2D rotation matrix $R(\theta)$ and its time derivative $\dot{R}(\theta,\dot{\theta})$ to calculate the new coordinates and velocities of the grid nodes.
Figure 5: Flow chart of the fluid–structure coupling. Within every physical time step, internal iterations between the aerodynamics and rigid body dynamics solver are made to ensure a strong coupling and high–fidelity FSI computations.
3.4 Sensitivity analysis
In this section, a comprehensive sensitivity analysis of the employed numerical parameters is presented, aiming at minimizing numerical uncertainties and ensuring the robustness and validity of the predicted results. The core objectives of this section involve the rigorous investigation of grid and temporal discretisation effects. This study focuses specifically on a highly challenging LAO case at V=24m/s airspeed from the dimitriadis2009bifurcation experimental campaign, as this regime imposes the most stringent numerical requirements.
3.4.1 Grid Dependency Analysis
The initial focus of the sensitivity analysis is the refinement of the computational grid. An O–type mesh is utilised, extending to a far–field radius of one hundred chord lengths ($100c$) (see Figure 6). A refined near–wake region is employed to properly resolve the emitted vortices, defined by a characteristic length of $\Delta w=5*10^{-2}c$. This region extends $5c$ in the chordwise direction ($0.5c$ upstream and $4.5c$ downstream) and $4c$ in the normal direction ($2c$ towards the pressure and $2c$ towards the suction side). The airfoil surface is discretised in $N_{c}=1148$ cells along its perimeter, varying from a minimum spacing of $\Delta X_{min}=10^{-4}c$ at the leading and trailing edges to a maximum of $\Delta X_{max}=2*10^{-3}c$ close to the maximum thickness region, with an increase ratio of $1.1$. A structured mesh zone expands radially from the airfoil wall to ensure the correct prediction of the boundary layer development. The height of the first cell is $h_{1}=10^{-5}c$ (thereby ensuring a normalized wall distance of $y^{+}\leq 1$) and the total number of cells in the radial direction is $35$ with an increase ratio of $1.1$. For efficiency, an unstructured meshing technique is employed for the remainder of the domain, with a growth rate of $1.1$, resulting in a total number of approximately $105000$ cells.
Figure 6: Overview of the O–type computational grid employed for the aeroelastic simulations. The domain extends to a far–field radius of $100c$ and totals approximately $105000$ cells. It consists of $N_{c}=1148$ cells along the airfoil perimeter, with a minimum discretisation length of $\Delta X_{min}=10^{-4}c$ at the leading and trailing edge and a growth rate of $1.1$ towards the maximum thickness region where a maximum length of $\Delta X_{max}=2*10^{-4}c$ is maintained. The boundary layer region comprises of $35$ cells with a first cell height of $h_{1}=10^{-5}c$ (ensuring $y^{+}\leq 1$) and a growth ratio of $1.1$, whereas a refined near–wake region with a characteristic length of $\Delta w=5*10^{-2c}$ is used to capture the dynamic vortex shedding.
A grid dependency analysis has been performed across four levels of refinement (Coarse, Medium, Fine and Extra Fine) to assess the variation in the predicted LCO characteristics (Amplitude and Period) listed in Table 1. All tested grids maintain a fixed minimum spacing of $\Delta X_{min}=10^{-4}c$ at the leading and trailing edges but differ in the maximum spacing ($\Delta X_{max}$), the number of cells defining the airfoil ($N_{c}$), and the characteristic length in the near–wake region ($\Delta w$). The Coarse grid ($\approx 60000$ cells) uses $N_{c}=658$ cells to describe the airfoil shape, with a maximum spacing of $\Delta X_{max}=4*10^{-3}c$ and a characteristic length of $\Delta w=10^{-1}c$ in the near wake region. It demonstrates significant inaccuracy, underestimating the Pitch Amplitude by approximately $9.4\%$ and overestimating the Period by approximately $19.4\%$ relative to the Extra Fine mesh ($\approx 440,000$ cells) that uses $N_{c}=2130$ cells in the airfoil wall, with a maximum spacing of $\Delta X_{max}=10^{-3}c$ and a characteristic length of $\Delta w=1.25*10^{-2}c$ in the near wake region. The Medium grid ($\approx 105000$ cells) uses $N_{c}=1148$ cells to describe the airfoil shape, with a maximum spacing of $\Delta X_{max}=2*10^{-3}c$ and a characteristic length of $\Delta w=5*10^{-2}c$ in the near wake region and accomplishes substantially improved convergence, with variations of only approximately $0.1\%$ in Amplitude and $3.5\%$ in Period. The Fine mesh ($\approx 215000$ cells) uses $N_{c}=2130$ cells in the airfoil wall, with a maximum spacing of $\Delta X_{max}=10^{-3}c$ and a characteristic length of $\Delta w=2.5*10^{-2}c$ in the near wake region and predicts similar deviations of approximately $0.7\%$ and $4.5\%$ in Amplitude and Period respectively. Consequently, the Medium mesh has been selected as the optimal compromise, providing accurate results at a manageable computational cost.
Table 1: Variation of Amplitude $^{\circ}$] and Period [sec] with grid refinement. Reference Amplitude and Period correspond to Extra Fine mesh comprising of $\approx 440000$ cells. The Medium mesh ($\approx 105000$ cells) provides the best compromise between accuracy and computational cost.
#cells\#\>cells Amplitude $^{\circ}$] Period [sec]
$\approx 60k$ $21.31^{\circ}$ $(-9.4\%)$ $1.3175$ $(+19.4\%)$
$\approx 105k$ $23.54^{\circ}$ $(+0.1\%)$ $1.1423$ $(+3.5\%)$
$\approx 215k$ $23.36^{\circ}$ $(-0.7\%)$ $1.1535$ $(+4.5\%)$
$\approx 440k$ $23.52^{\circ}$ ()(-) $1.1037$ ()(-)
A specific analysis on the minimum spacing, $\Delta X_{min}$, has been also conducted, highlighting its critical influence on aeroelastic predictions. The effect of $\Delta X_{min}$ has been previously overlooked in the respective literature, where grid dependency analyses are usually restricted solely to computing static polars watrin2012computational; yabili2012unsteady; poirel2011computational; abbas2021aeroelastic; li2024mechanism. Table 2 provides a clear overview of the effect of $\Delta X_{min}$ on the Amplitude and Period of the predicted LCO, clearly indicating that $\Delta X_{min}=10^{-4}c$ yields effectively in a grid–independent solution. On the other hand, when the coarse leading and trailing edge discretisation ($\Delta X_{min}=10^{-3}c$) is employed, Pitch Amplitude is overestimated by approximately $6.7\%$ and the LCO Period is underestimated by approximately $3.7\%$. Figure 7 further illustrates this by showing the time history predicted when different values of $\Delta X_{min}$ are employed, where it is shown that the size of $\Delta X_{min}$ clearly affects the Amplitude and Period of the predicted pitching motion. This sensitivity stems from the influence of the grid resolution, particularly at the leading edge, on the precise development and subsequent detachment dynamics of the Leading Edge Vortex (LEV). Crucially, the flow field and pressure distribution over the airfoil presented in Figure 8 are captured at the time instant denoted by the red circle in Figure 7. At this specific time, the differences observed in the aerodynamic field and thus the resulting aerodynamic moment are mainly attributed to the $\Delta X_{min}$ selection, as the airfoil exhibits similar pitch angle (see Figure 7(a)) and angular velocity (see Figure 7(b)) across the tested discretisations up to this point. When the leading edge discretisation is coarser, the LEV finds it challenging to detach, remaining closer to the airfoil (see Figure 8(a)). This proximity leads to higher local velocities and thus increased suction aft from the pitching axis (see Figure 8(b)). Ultimately, this leads to an increased excitation aerodynamic moment at this time instant that later results in higher pitch angles (see the different peak values predicted in 7 right after the red dot). Conversely, finer discretisation facilitates smoother LEV shedding and convection, moving the vortex farther from the airfoil and thereby generating an increased restoring aerodynamic moment that drives the airfoil to lower pitch peak values after the red dot time instant. These complex, non–trivial dependencies, visually supported by the flow field structures presented in Figure 8, demonstrate conclusively that grid dependency analyses on Stall Flutter instabilities must be extended beyond standard criteria, focusing specifically on how dynamic flow phenomena like vortex shedding are affected, rather than merely relying on static airfoil polar predictions.
Table 2: Variation of Amplitude $^{\circ}$] and Period [sec] with decreasing Leading Edge and Trailing Edge spacing Xmin[m]Xmin\>[m], defined with respect to airfoil chord $c$. Reference Amplitude and Period correspond to minimum value $Xmin=10^{-}5c$. $Xmin=10^{-}4c$ yields in a grid–independent solution.
Xmin[]Xmin\>[-] Amplitude $^{\circ}$] Period [sec]
$10^{-}3c$ $25.16^{\circ}$ $(+6.7\%)$ $1.0982$ $(-3.7\%)$
$10^{-}4c$ $23.54^{\circ}$ $(-0.2\%)$ $1.1423$ $(+0.1\%)$
$10^{-}5c$ $23.58^{\circ}$ ()(-) $1.1408$ ()(-)
(a) Pitch time history
(b) Pitch rate time history
Figure 7: Impact of minimum spacing ($\Delta X_{min}$) at the leading and trailing edges on pitch angle (a) and pitch rate (b) time history. All three simulations have a similar pitch angle and pitch rate at this instant. Consequently, the deviations in the predicted pitch peak after the red dot can be attributed to different aerodynamic excitation at this time instant.
(a) Vorticity controur
(b) $C_{p}$ distribution
Figure 8: Vorticity contour (a) and pressure coefficient distribution (b) for different leading edge and trailing edge discretisation lengths generated at the time instant (noted with a red dot in Figure 7). The influence of leading edge discretisation on leading edge vortex dynamics is highlighted, indicating that coarser discretisation hinders the detachment of the vortex, causing it to remain closer to the airfoil surface and resulting in an artificially increased excitation aerodynamic moment and thus larger predicted LCO amplitude.
3.4.2 Time–step Dependency Analysis
Table 3 shows how the selected non–dimensional time–step $\Delta\tau=\frac{\Delta t\>V}{c}$ affects the aeroelastic response of the system. Except from the value of $\Delta\tau=0.02$ that results in ever–increasing amplitude, a monotonic convergence is observed, with $\Delta\tau=0.005$ yielding a time–step independent solution. This non–dimensional time-step value ensures that at least $14400$ steps per oscillation period will be utilised to simulate the aeroelastic response of the system. Based on this outcome, the time–step value $\Delta t$ turns out to be yet another numerical parameter that shows increased sensitivity when investigated within an aeroelastic context. This is clearly shown in Figure , where the impact of the time–step value in the predicted lift coefficient variation is depicted for different oscillating conditions. In Figure 9(a) it is shown that a $\Delta t$ corresponding to only $3600$ steps per period is found to be sufficient for a time–step independent prediction of a dynamically pitching airfoil, whereas more than $14400$ steps per period are necessary for the aeroelastic case shown in Figure 9(b). This lenient requirement resulting from dynamic cases of rigid airfoils undergoing imposed pitching motion is considered as standard practice for time–step independence analyses found in literature. Nevertheless, it is clearly shown in this study that the time–step value required to correctly resolve the aeroelastic response of an airfoil undergoing stall flutter oscillations is significantly smaller.
Table 3: Variation of Amplitude $^{\circ}$] and Period [sec] with decreasing non–dimensional time–step value $\Delta\tau=\frac{\Delta t\>V}{c}[-]$. Reference Amplitude and Period correspond to finest value $\Delta\tau=0.0025$. $\Delta\tau=0.005$ yields in a time–step independent solution.
$\Delta\tau\>[-]$ Amplitude $^{\circ}$] Period [sec]
$0.04$ $34.14^{\circ}$ $(+45\%)$ $0.7853$ $(-31.2\%)$
$0.02$ $\infty$ $\infty$
$0.01$ $25.80^{\circ}$ $(+9.6\%)$ $0.971$ $(-14.8\%)$
$0.005$ $23.54^{\circ}$ $(+0\%)$ $1.1423$ $(+0.1\%)$
$0.0025$ $23.54^{\circ}$ ()(-) $1.1412$ ()(-)
(a) Dynamically pitching NACA0012 (imposed motion) at $a_{0}=10^{\circ}$, $a_{1}=15^{\circ}$ and $k=0.1$, at $Re=135000$ and $Ma=0.103$. Experimental data from lee2004investigation and numerical predictions from liggett2012temporal.
(b) Predictions by MaPFlow for an aeroelastically pitching NACA0012 in symmetric LAO at $Re=540000$ and $Ma=0.079$, firstly observed in dimitriadis2009bifurcation experimental campaign.
Figure 9: Lift coefficient variation with pitch angle in dynamic (a) and aeroelastic conditions (b) for a pitching NACA0012. The time–step value required to correctly resolve the aeroelastic response of a pitching airfoil (b) is significantly smaller compared to the one needed to accurately predict the aerodynamic loading of a rigid airfoil undergoing imposed pitching motion (a).
4 Results
In this section, the numerical results are presented by applying the validated methodology and numerical setup presented in Section 3. The investigation focuses on the two distinct Stall Flutter regimes: SAO in the transitional Reynolds number range and LAO in the moderate Reynolds number regime.
4.1 Small Amplitude Oscillations
In this sub–section, the aeroelastic response of a system that results in SAO is examined. This phenomenon has been observed by poirel2008self in a low and constrained transitional Reynolds number range (from $45000$ to $130000$). In poirel2011computational it is shown that the flow separation close to the trailing edge plays a determinant role in the onset of the fluttering motion, while the subsequent presence of a LSB stabilizes this motion into a LCO. MaPFlow predictions, illustrated in Figure 10, confirm this mechanism, showing that the emergence of trailing edge vortices at time–stand "11" drives the system towards an instability onset, which is subsequently stabilized by the formation of larger detached vortices near the trailing edge (time–stands "22" and "33"). Hence, the employment of transition modelling is considered essential to correctly capture the initial trailing edge separation, consequently leading to the SAO regime. Neglecting this transition model in the numerical simulations results in the complete suppression of the LCO.
Figure 11 presents a comprehensive comparison of the resulting LCO characteristics, specifically amplitude and frequency, for various Reynolds numbers in the transitional range. Experimental measurements are juxtaposed against numerical predictions obtained using various turbulence modelling options. It is evident that Laminar and the standard Fully Turbulent simulations deviate significantly from the experimentally measured trend in amplitude (see Figure 11(a)). However, the non–monotonic trend of increasing amplitude at lower Reynolds numbers ($Re\leq 85400$) followed by a decrease as the Reynolds number further increases, is successfully captured by the Low Reynolds Correction and the Transitional simulations. Nevertheless, a persistent over–prediction of amplitudes in the higher Reynolds range is shown, where computational models predictions agree well with each other. This amplitude overestimation directly correlates with the prediction of lower oscillation frequencies by the present study simulations in the same range (see Figure 11(b)). However, the overall trend of increasing oscillation frequency with increasing Reynolds number is fairly predicted.
Figure 10: Pitch response time history, vorticity contours, friction and pressure coefficient distribution for a SAO case at a Reynolds number of $Re=1.17*10^{5}$, firstly observed in poirel2008self experimental campaign. Three critical time instants have been makred in the time history plot, namely: "11" where the emergence of trailing edge vortices leads to an instability onset that drives the system to oscillations, "22 and "33" where the subsequent formation of larger detached vortices stabilizes the system to LCO. The corresponding friction and pressure coefficient distributions imprint the flow dynamics associated with the vortices formation. Red dots in friction coefficient plots indicate the laminar separation positions where LSBs form.
(a) Pitch Amplitude
(b) Pitch Frequency
[…]
4.2 Large Amplitude Oscillations
[…]
The structural damping in the experimental setup (primarily due to friction in the pitch springs) was found to exhibit highly non–linear characteristics. Experimental observations, specifically mentioned in Section , indicated that the critical pitch damping coefficient $\zeta$ demonstrated pronounced non–linearity primarily near small pitch angles (around the zero position) and transitioned to a linear increase at higher pitch angles. This behavior necessitates the implementation of a non–linear damping model, typically structured as a function of the pitch angle $\zeta(\theta)$, as describe by Equation (9) and illustrated in Figure 12. Although the damping was noted to increase linearly at larger pitch angles, the component representing this linear increase has been entirely omitted in the final model. This simplification is justified by the fact that the corresponding angular velocity $\dot{\theta}$ is typically small at large angles, meaning that the contribution of this linear term to the overall off–wind response is insignificant. Omitting this term consequently reduces the difficulty of tuning the model by removing an additional unknown variable. Figure 13 presents the successful validation of the chosen non–linear damping function of Equation (9) against experimental free–decay tests. The comparison confirms that this model accurately captures the dynamic response of the system in the absence of significant aerodynamic excitation, illustrating the characteristic rapid decay of oscillations observed experimentally.
$\begin{split}\zeta(\theta)&=0.52e^{-0.03\theta^{2}}\\ C(\theta)&=2\zeta(\theta)\sqrt{IK}\end{split}$ (9)
Figure 12: Non–linear critical damping ratio $\zeta$ with respect to pitch angle $\theta$ (in degrees) employed in LAO simulations of dimitriadis2009bifurcation cases.
(a) V=0m/sV=0\;m/s
(b) V=2.5m/sV=2.5\;m/s
(c) V=3.1m/sV=3.1\;m/s
Figure 13: Numerical prediction of the free–decay tests performed at 0 (a), $2.5$ (b) and $3.1$ (c) $m/s$ airflow speed. The comparison against experimental measurements confirms that the non–linear damping model described in Equation (9) accurately captures the dynamic response of the system in the absence of significant aerodynamic excitation.
Figure 14 illustrates the transition from static stability to LCO for increasing values airflow velocity. When the aeroelastic system is exposed to low airflow velocity, such as V=18m/sV=18\;m/s, the airfoil deflects to its respective static equilibrium angle by following the typical response of a $2nd$ order dynamic system. Although the simulations start with the airfoil at $0^{\circ}$ pitch angle, where negligible aerodynamic lift or moment is theoretically expected, even a minor asymmetry inherent in the computational grid can induce an initial asymmetric loading over the airfoil. Consequently, the airfoil deflects to a non–zero pitch angle, which subsequently establishes a non–zero lift calculation. Provided that the distance between the pitching axis and the quarter–chord is sufficiently large, and considering that the angular velocity of the resulting response is small, the aerodynamic moment around the pitching axis will overcome the restoring damping moment, causing the airfoil to deflect to progressively larger pitch angles, dynamically heading to the static equilibrium pitch angle. (In order to accelerate this initial phase of the simulations, the initial pitch angle of the airfoils analysed herein is set to $0.2^{\circ}$ nose–up.) As velocity increases to V=18.4m/sV=18.4\;m/s, the balance between aerodynamic energy input and system damping becomes critical, resulting in marginal stabilisation to the static equilibrium pitch angle. A further increase to V=18.6m/sV=18.6\;m/s leads to the observed numerical critical velocity, where self–starting and sustained oscillations are predicted. At this critical velocity, the simulation initially exhibits a bifurcation from the symmetric to the asymmetric aeroelastic mode, subsequently resulting in small amplitude asymmetric oscillations around the static equilibrium angle. For velocities exceeding $V_{c,num}=18.6\;m/s$, symmetric oscillations of significantly larger amplitude around the zero mean pitch angle are predicted. The phase plane plot presented in Figure 15 clearly depicts these trajectories, confirming the shift from stable static equilibrium (V<18.6m/sV<18.6\;m/s) to stable motion of asymmetric SAO (V=18.6m/sV=18.6\;m/s) or symmetric LAO (V>18.6m/sV>18.6\;m/s).
Figure 14: Pitch angle time history for various airflow velocities demonstrating the transition from static stability to LCO. Cases shown include stable decay to the static equilibrium position (V=18m/sV=18\;m/s), marginally stable convergence (V=18.4m/sV=18.4\;m/s) and the onset of self–started and sustained oscillations at the numerical critical velocity ($V_{c,num}=18.6\;m/s$) where asymmetric oscillations around the respective static equilibrium angle are predicted along with a bifurcation from the symmetric to the asymmetric aeroelastic mode. For higher airflow velocities, only symmetric oscillations of significantly larger amplitude around the zero mean pitch angle are predicted.
Figure 15: Phase plane trajectories ($\theta$ versus $\dot{\theta}$) for various airflow velocities. Trajectories depict stable convergence to the static equilibrium point (V=18m/sV=18\;m/s), marginally stable convergence (V=18.4m/sV=18.4\;m/s), small amplitude asymmetric oscillations near the critical velocity (V=18.6m/sV=18.6\;m/s), and large amplitude symmetric oscillations around zero mean pitch angle (V=18.8m/sV=18.8\;m/s).
Figure 16 compares the computationally predicted LCO characteristics (Amplitude and Period variation with respect to the airflow speed) against the experimental measurements. The distinct representation of "Negative pitch centering" and "Positive pitch centering" measured data points originates in a small asymmetry in the experimental wing that drove the zero lift angle of attack to lie between $1^{\circ}$ and $2^{\circ}$, the positive stall angle to be at $13^{\circ}$, the negative stall angle at $10^{\circ}$ and the pitching moment around the quarter–chord to be small, but non–zero at small angles of attack. Hence, in an early phase of the experimental campaign, the wing was centered such that the rest pitch angle was slightly negative, leading to the occurrence of only negative angle asymmetric LCO. The dynamic experiments were then repeated with an alternative centering for the torsional springs that led to a slightly positive equilibrium pitch angle, which in turn led to the observation of positive pitch angle asymmetric LCO. For this reason, only the "Positive pitch centering" symbol is encountered in the positive minimum values of large airflow speeds and only the "Negative pitch centering" symbol in the negative minimum values. Nevertheless, the maximum values agree well with each other on both experimental configurations. On the other hand, the numerical configuration is purely symmetric, yielding in the same absolute value predictions for both positive and negative pitch angles. Yet, for airflow velocities greater than the critical one, only symmetric oscillations around zero mean value are predicted by the numerical model.
Firstly, the numerical predictions clearly over–predict the critical velocity ($V_{c,num}=18.6\;m/s$), which is measured at $V_{c,exp}\approx 13\;m/s$ (see Figure 16(a)). Furthermore, the predicted pitch amplitude at the onset ($A_{c,num}\approx 14.3^{\circ}$) is substantially lower than the measured amplitude at the experimental critical speed ($A_{c,exp}\approx 34.1^{\circ}$). This under–prediction continues systematically across the measured velocity range, where the predicted plateau amplitude ($\approx 29^{\circ}$) remains approximately $13^{\circ}$ lower than the measured plateau ($\approx 42^{\circ}$). This amplitude discrepancy explains the larger oscillation period observed in Figure 16(b) and is directly linked to the lower aerodynamic excitation peaks ($C_{M}$ and $C_{L}$) predicted in the numerical simulations, as illustrated in Figures 16(c) and 16(d) respectively. It needs to be highlighted that, following the approach of li2007experimental, high frequency peaks are neglected in the process of identifying the maximum and minimum values of $C_{M}$ and $C_{L}$. Despite these differences in magnitude, the simulation successfully captures the overall qualitative trend of increasing amplitude followed by a plateau, and the corresponding reduction in period with increasing airflow velocity.
Furthermore, the sectional pitching moment and lift hysteresis curves (Figures 17 and 18) demonstrate strong qualitative agreement with the experimental data. The same outcome can be made also for Figure 19, where only the asymmetric part of the loop has been measured. Specifically, MaPFlow predicts quite well the minimum values of moment, along with the folding of the curve when the airfoil approaches the minimum pitch value. The disagreement in the respective maximum values of the pitching moment at positive pitch angles originates in the asymmetric shape of the experimental wing. Furthermore, the almost constant values of moment during the upstroke motion at negative pitch angles and the downstroke motion at positive angles is also successfully predicted. The high frequency ripples detected in the predicted moment around these almost constant values are not evident in measurements and are related to the break down of the large LEVs shown in time instants "44" and "88" of Figure 20, to the smaller structures shown in time instants "55" and "99". MaPFlow is also able to predict the extension of the linear phase in post–static–stall angles of attack, by successfully resolving the delay of the boundary layer separation from the airfoil surface during upstroke at positive pitch angles (see time instant "11" in Figure 20) and downstroke at negative angles (see time instant "66" in Figure 20). The closed loops predicted around $0^{\circ}$ pitch angle (they are more easily identified in moment curves) are also evident in measurements, but slightly shifted due to the experimental wing asymmetry. For the same reason, the measured slope of the linear part of the lift curves at positive pitch angles is slightly lower than that of the negative angles.
In summary, MaPFlow successfully predicts the fundamental dynamics and key features of the aeroelastic system analysed in this Section. Both symmetric pitch oscillations around zero mean value and asymmetric ones around the static equilibrium angle have been successfully predicted. However, the asymmetric oscillations, along with a bifurcation from the symmetric to the asymmetric aeroelastic mode, are encountered by the simulation only close to the critical velocity, where the amplitude of the oscillations is relatively small. The critical velocity is over–predicted compared to the experimentally measured one, which, however,has been found to be a common tendency among various computational tools’ predictions also in the SAO Section 4.1 analysed before. Furthermore, the systematic pitch amplitude under–prediction in all the examined airflow velocities agrees well with the respective larger oscillation period predictions and is attributed to the corresponding lower sectional moment and lift coefficients predicted in the symmetric oscillations velocities (up to 25m/s25\;m/s). Nevertheless, apart from the narrower pitch angle range, the sectional aerodynamic loads predictions with respect to pitch angle agree qualitatively well with the wind tunnel measurements throughout the computed velocity range.
(a) Pitch
(b) Period
(c) $C_{M}$
(d) $C_{L}$
Figure 16: Comparison of computationally predicted LCO characteristics against experimental measurements from dimitriadis2009bifurcation. Figures show the variation with airflow velocity of Pitch Amplitude (a), Period (b), maximum Pitching Moment Coefficient (c) and maximum Lift Coefficient (d). Systematic discrepancies include the over–prediction of the critical velocity ($V_{crit}$) and the under–prediction of the LCO amplitudes across the velocity range, along with the respective larger oscillation Period.
(a) $C_{M}$
(b) $C_{L}$
Figure 17: Sectional Pitching Moment Coefficient $C_{M}$ (a) and Lift Coefficient $C_{L}$ (b) variation with respect to pitch angle at V=23m/sV=23\;m/s airspeed. The hysteresis loops demonstrate strong qualitative agreement with wind tunnel measurements, successfully capturing non–linear folding behavior near stall and the extension of the linear phase due to dynamic flow effects.
(a) $C_{M}$
(b) $C_{L}$
Figure 18: Sectional Pitching Moment Coefficient $C_{M}$ (a) and Lift Coefficient $C_{L}$ (b) variation with respect to pitch angle at V=24m/sV=24\;m/s airspeed. The hysteresis loops demonstrate strong qualitative agreement with wind tunnel measurements, successfully capturing non–linear folding behavior near stall and the extension of the linear phase due to dynamic flow effects.
(a) Sectional Moment
(b) Sectional Lift
Figure 19: Sectional Pitching Moment (a) and Lift (b) variation with respect to pitch angle at V=27m/sV=27\;m/s airspeed. The hysteresis loops demonstrate strong qualitative agreement with wind tunnel measurements, successfully capturing non–linear folding behavior near stall and the extension of the linear phase due to dynamic flow effects, despite the observation of only the asymmetric part of the loop in the wind tunnel experiment.
Figure 20: Prediction of sectional Pitching Moment Coefficient $C_{M}$ and Lift Coefficient $C_{L}$ variation with respect to pitch angle during one symmetric LCO cycle at V=24m/sV=24\;m/s (top) and instantaneous vorticity contours (bottom). The sequence illustrates the dynamic stall mechanism, including the generation of a strong LEV during upstroke at positive pitch angles (time instants "11" to "33"), the shedding and breake–down of the LEV into smaller vorticity structures during the downstroke at positive pitch angles (time instants "44" to "55") and the respective phenomenon at negative pitch angles (downstroke in time instants "66" to "77" and upstroke in time instants "88" to "99"). A strong fluid–structure interaction dynamics is effectively captured by the simulation.
5 Conclusion
The objective of this paper is to perform a comprehensive numerical study investigating the proper prediction of small and large amplitude LCO arising from Stall Flutter instabilities. Computational results are compared against experimental measurements from two distinct campaigns performed at low and moderate Reynolds numbers and numerical predictions by other computational tools of similar fidelity wherever available. The aeroelastic response of the systems was computed using MaPFlow, an in–house compressible URANS solver that employs Low Mach Preconditioning for accurate analysis of low Mach number flow regions, coupled strongly with a rigid body dynamics solver.
The investigation includes a detailed sensitivity analysis of numerical parameters, focusing on grid and temporal discretization, which are critical for accurate FSI predictions. For the grid dependency, a grid consisting of approximately $105000$ cells ($N_{c}=1148$ airfoil cells) is selected as the optimal trade–off between accuracy and computational cost. A minimum spacing $\Delta X_{min}=10^{-4}c$ at the leading and trailing edge has been found to be necessary to achieve a grid–independent prediction for LCO amplitude and period. In the temporal analysis, a non–dimensional time–step value of $\Delta\tau=\dfrac{\Delta tV}{c}=0.005$, ensuring a minimum of $14400$ steps per period, is required for a time–step independent solution in the aeroelastic context. This is a significantly stricter requirement than that found sufficient for dynamically pitching rigid airfoils, where $1800$ steps per period provide a qualitatively correct numerical prediction.
The computational framework presented herein is able to fairly predict both the aerodynamic loads and the dynamic response of the aeroelastic systems analyzed. Both symmetric pitch oscillations around a zero mean value and asymmetric oscillations around the static equilibrium angle are successfully predicted. Asymmetric oscillations, along with a bifurcation from the symmetric to the asymmetric aeroelastic mode, are observed only near the critical velocity, where the oscillation amplitude is relatively small.
[…]
For LAO, the critical velocity is again over–predicted by MaPFlow. Furthermore, the predicted pitch amplitudes are generally lower than the measured values (predicted plateau at $\approx 29^{\circ}$ versus measured plateau at $\approx 42^{\circ}$), resulting in greater oscillation periods. This under–prediction of amplitude is attributed to the lower aerodynamic loads (excitation) peaks reported in the symmetric oscillations region (up to 25m/s25\;m/s). Despite these amplitude discrepancies, the sectiotanl pitching moment and lift predictions show qualitatively good agreement with wind tunnel measurements.
In conclusion, the aeroelastic code provides comparable results to those produced by standard CFD solvers of similar accuracy. While the qualitative agreement for complex flow physics like sectional loads and bifurcation behavior is strong.
Acknowledgments
This project is carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union – NextGenerationEU (H.F.R.I. Project Number: 016749)
The authors gratefully acknowledge the HPC RIVR consortium [(www.hpc-rivr.si)](https://www.hpc-rivr.si) and EuroHPC JU [(eurohpc-ju.europa.eu)](https://eurohpc-ju.europa.eu/) for funding this research by providing computing resources of the HPC system Vega at the Institute of Information Science [(www.izum.si)](https://www.izum.si/en/home/).
The computational grids were generated using the ANSA CAE pre-processor of “BETA_CAE Systems S.A”.